# GSAlign: Geometric and Semantic Alignment Network for Aerial-Ground Person Re-Identification

### v1

@@ Abstract
Aerial-Ground person re-identification (AG-ReID) is an emerging yet challenging task that aims to match pedestrian images captured from drastically different viewpoints, typically from unmanned aerial vehicles (UAVs) and ground-based surveillance cameras. The task poses significant challenges due to extreme viewpoint discrepancies, occlusions, and domain gaps between aerial and ground imagery. While prior works have made progress by learning cross-view representations, they remain limited in handling severe pose variations and spatial misalignment. To address these issues, we propose a Geometric and Semantic Alignment Network (GSAlign) tailored for AG-ReID. GSAlign introduces two key components to jointly tackle geometric distortion and semantic misalignment in aerial-ground matching: a Learnable Thin Plate Spline (LTPS) Module and a Dynamic Alignment Module (DAM). The LTPS module adaptively warps pedestrian features based on a set of learned keypoints, effectively compensating for geometric variations caused by extreme viewpoint changes. In parallel, the DAM estimates visibility-aware representation masks that highlight visible body regions at the semantic level, thereby alleviating the negative impact of occlusions and partial observations in cross-view correspondence. A comprehensive evaluation on CARGO with four matching protocols demonstrates the effectiveness of GSAlign, achieving significant improvements of +18.8% in mAP and +16.8% in Rank-1 accuracy over previous state-of-the-art methods on the aerial-ground setting. The code is available at: https://github.com/stone96123/GSAlign.
@@ 1 Introduction
Person re-identification (ReID), aiming to address the problem of matching people over a distributed set of nonoverlapping cameras, has attracted intensive attention in the last few years due to its wide applications in surveillance systems. Although traditional person re-identification (ReID) has achieved remarkable progress [1, 2, 3, 4, 5], there is growing interest in aerial-ground ReID (AG-ReID) settings, driven by the increasing deployment of unmanned aerial vehicles (UAVs) and low-altitude platforms in surveillance applications [6, 7, 8, 9, 10]. This, in practice, puts the Re-ID problem in an aerial-ground setting and requires the approaches to properly handle both the significant geometric view-variation and semantic misalignment.
Despite its potential, AG-ReID remains a highly underexplored and technically challenging problem. Unlike conventional ReID tasks [11, 12, 13, 14, 15], AG-ReID suffers from extreme viewpoint disparities, where aerial images exhibit severe top-down perspectives while ground views contain mostly frontal or profile views. These differences cause significant geometric distortions and drastic appearance changes, which conventional ReID models that are primarily trained on ground-view data are ill-equipped to handle. In addition, AG-ReID often involves frequent occlusions and unbalanced visibility across views, further complicating cross-view matching.
To mitigate these challenges, several recent studies have attempted to adapt existing ReID techniques to the aerial-ground scenario [16, 8]. These methods seek to learn modality-specific representations or map features into a shared embedding space using metric learning or adversarial objectives. Others employ cross-view decomposition strategies or design view-invariant constraints to bridge the modality gap. However, these methods tend to focus on global alignment and overlook two critical issues: (1) the severe geometric distortion induced by cross-view perspectives, and (2) semantic misalignment caused by partial occlusions and varying visible body regions. These limitations lead to suboptimal performance, especially under large viewpoint gaps.
@@ 2 Related Work
Building upon well-established ground surveillance infrastructure, advanced person re-identification algorithms have made significant progress across a range of scenarios, including general scenarios [17, 18, 1, 4], occluded scenarios [19, 12, 20, 21], cross-modal scenarios [22, 23, 24, 25], multi-spectral scenarios [26, 27, 28], and cross-spectral scenarios [29, 30, 31, 32, 33].
However, with the proliferation of unmanned aerial vehicles (UAVs) and low-altitude platforms augmenting traditional surveillance systems, person ReID research is increasingly extending to aerial-view scenarios [34]. Initial efforts in this direction focused on aerial-aerial retrieval tasks using datasets such as UAV-Human [35] and PRAI-1581 [7]. More recently, attention has shifted towards connecting the aerial views with traditional ground systems. Aerial-Ground person ReID (AG-ReID), where the query and gallery images originate from fundamentally different viewpoints, typically aerial and ground-level perspectives. This setup introduces severe spatial and semantic discrepancies, often leading to substantial degradation in identity-preserving cues such as pose, silhouette, and clothing texture [36]. Several approaches have been proposed to mitigate these challenges. Nguyen et al. [6] proposed a dual-stream structure guided by attributes to enhance semantic disentanglement, later extended to a three-stream model with modality-aware supervision [16]. Zhang et al. [8] developed the View-Decoupled Transformer (VDT), which explicitly models and separates viewpoint-specific and viewpoint-invariant features. Despite these advances, existing AG-ReID methods still face difficulties in handling extreme spatial distortions and in leveraging visible semantic regions effectively. To overcome these limitations, we propose GSAlign, a novel framework that performs instance-adaptive geometric transformation and visibility-aware semantic alignment, promoting robust and spatially consistent feature learning across aerial and ground domains.
@@ 3.1 Overview
The overall framework of the Geometric and Semantic Alignment Network (GSAlign), illustrated in Fig. 2, is built upon a ViT-Base backbone and designed to address feature misalignment under complex multi-view conditions. Inspired by the success of the View-Decoupled Transformer (VDT) [37], we adopt the VDT architecture as the foundation of GSAlign, which effectively captures both global semantics and fine-grained structural cues. Following its design, we introduce an additional view token to model viewpoint-specific information. The token sequence during training is defined as:
@@ 3.3 Dynamic Alignment Module
In person re-identification, training images often contain occlusions, background noise, or identity-irrelevant regions. Directly encoding features from the entire image may lead the model to learn redundant or non-discriminative representations. To address this issue, we explore whether the visibility patterns of a given input image can be used to dynamically modulate the feature representations of other samples. Inspired by DPM [38], we propose the Dynamic Alignment Module (DAM), which leverages input-guided masks to improve semantic alignment. Specifically, DAM treats the mean feature representation of images with the same identity as its ideal prototype, assuming that noise or occlusion in different samples manifests as partial loss of this prototype’s semantic structure. Based on this formulation, DAM introduces a dynamic channel-wise mask generator based on the input and modulates prototypes by referencing the input-aware mask. This allows the model to suppress irrelevant or noisy dimensions and emphasize input-relevant subspaces in a content-aware manner.
@@ 4.1 Implementation
Datasets and Evaluation Protocols. We conduct experiments on the CARGO dataset [8], a large-scale benchmark specifically designed for aerial-ground person re-identification (AG-ReID). CARGO contains 108,563 images of 5,000 synthetic identities captured by 13 cameras (5 aerial and 8 ground), with diverse conditions including extreme viewpoint changes, resolution variations, illumination shifts, and occlusions. The dataset is constructed in a simulated urban environment using MakeHuman for identity modeling and Unity3D for camera deployment.
Following the standard protocol, we use 51,451 images with 2500 identities for training and 51,024 images with the remaining 2,500 identities for testing. The testing phase contains four evaluation protocols: ALL, ground to ground( GG), aerial to aerial ( AA), and aerial to ground (AG), where each protocol focuses on a specific matching scenario. To further verify the generalization ability of GSAlign in real-world scenarios, we also conduct comparative experiments on the AG-ReID [6] and AG-ReID v2 [16] datasets, which contain real aerial and ground imagery captured in outdoor environments. We use Cumulative Matching Characteristics (CMC) at Rank-1 accuracy, mean Average Precision (mAP), and mean Inverse Negative Penalty (mINP) [39] as evaluation metrics.
Implementation Details. All implementation settings strictly follow the protocol of VDT [8] for fair comparison. We adopt ViT-Base as the backbone, initialized with ImageNet-21K pretraining. Input images are resized to , and horizontal flipping is applied during training for data augmentation. The model is optimized using AdamW with weight decay of 0.05 and a cosine learning rate schedule. The initial learning rate is set to with linear warm-up for the first 20 epochs. is set to 0.1 during the training. We train for 120 epochs in total, using a batch size of 64.
@@ 4.2 Comparison with State-of-the-Arts
| SBS [40] | 50.32 | 43.09 | 29.76 | 72.31 | 62.99 | 48.24 | 67.50 | 49.73 | 29.32 | 31.25 | 29.00 | 18.71 |
| PCB [17] | 51.00 | 44.50 | 32.20 | 74.10 | 67.60 | 55.10 | 55.00 | 44.60 | 27.00 | 34.40 | 30.40 | 20.10 |
| BoT [41] | 54.81 | 46.49 | 32.40 | 77.68 | 66.47 | 51.34 | 65.00 | 49.79 | 29.82 | 36.25 | 32.56 | 21.46 |
| MGN [42] | 54.81 | 49.08 | 36.52 | 83.93 | 71.05 | 55.20 | 65.00 | 52.96 | 36.78 | 31.87 | 33.47 | 24.64 |
| VV [43, 44] | 45.83 | 38.84 | 39.57 | 72.31 | 62.99 | 48.24 | 67.50 | 49.73 | 29.32 | 31.25 | 29.00 | 18.71 |
| AGW [39] | 60.26 | 53.44 | 40.22 | 81.25 | 71.66 | 58.09 | 67.50 | 56.48 | 40.40 | 43.57 | 40.90 | 29.39 |
| BAU [45] | 45.20 | 38.40 | - | 61.60 | 51.20 | - | 50.00 | 42.60 | - | 40.40 | 36.70 | - |
| PAT [46] | 37.90 | 15.30 | - | 52.70 | 24.20 | - | 50.00 | 23.10 | - | 35.10 | 15.50 | - |
| DTST [47] | 64.42 | 55.73 | 41.92 | 78.57 | 72.40 | 62.10 | 80.00 | 63.31 | 44.67 | 50.53 | 43.49 | 29.46 |
| ViT [37] | 61.54 | 53.54 | 39.62 | 82.14 | 71.34 | 57.55 | 80.00 | 64.47 | 47.07 | 43.13 | 40.11 | 28.20 |
| VDT [8] | 64.10 | 55.20 | 41.13 | 82.14 | 71.59 | 58.39 | 82.50 | 66.83 | 50.22 | 48.12 | 42.76 | 29.95 |
@@
In this section, we compare our proposed GSAlign with a range of state-of-the-art person ReID methods, including both conventional CNN-based models (e.g., PCB [17], BoT [41], MGN [42]) and Transformer-based frameworks (e.g., ViT [37], VDT [8]). As shown in Table 1, GSAlign achieves the best overall performance across all four protocols. Specifically, under the most comprehensive ALL protocol, GSAlign achieves 65.06% Rank-1, 57.95% mAP, and 44.97% mINP, outperforming the strong baseline VDT [8] by +0.96%, +2.75%, and +3.84%, respectively. On the most challenging AG cross-view setting, our method surpasses all previous methods by a large margin, reaching 64.89% Rank-1, 61.55% mAP, and 52.81% mINP, significantly ahead of the second-best VDT by +16.77%, +18.79%, and +22.86%, respectively. Although GSAlign is not specifically designed for single-view matching protocols, it does not introduce any negative effects in such settings. On both AA and GG, GSAlign still achieves competitive performance. We attribute this to the fact that, in single-view scenarios, the features are already well aligned due to the limited viewpoint variation, thus requiring minimal additional geometric or semantic correction. In conclusion, these results clearly demonstrate that our proposed GSAlign not only achieves strong global retrieval accuracy but also exhibits exceptional robustness in the presence of extreme viewpoint changes, validating the effectiveness of our geometric and semantic alignment design.
@@
| ViT [37] | 81.47 | 72.61 | - | 82.85 | 73.39 | - |
| VDT [8] | 82.91 | 74.44 | 51.06 | 86.59 | 78.57 | 52.87 |
| GSAlign | 86.74 | 84.00 | 73.62 | 87.94 | 87.17 | 78.21 |
Furthermore, we evaluate the proposed GSAlign on the real-world AG-ReID [6] and AG-ReID v2 [16] datasets, and compare it against a broad range of state-of-the-art ReID approaches, as summarized in Table 2 and Table 3. On the AG-ReID dataset, GSAlign consistently achieves the best overall performance across both cross-view protocols. Specifically, under the AG setting, it attains 86.74% Rank-1, 84.00% mAP, and 73.62% mINP, outperforming the strong baseline VDT by large margins of +3.83%, +9.56%, and +22.56%, respectively. Under the inverse GA protocol, GSAlign maintains superior results with 87.94% Rank-1, 87.17% mAP, and 78.21% mINP, exceeding VDT by +1.35%, +8.60%, and +25.34%, respectively. These substantial improvements demonstrate the advantage of our geometric-semantic alignment mechanism in handling severe viewpoint discrepancies between aerial and ground cameras. For the AG-ReID v2 dataset, which introduces more realistic and diverse environmental conditions, GSAlign again achieves the best performance in most matching protocols. Under the comprehensive ALL protocol, GSAlign yields 91.47% Rank-1 and 89.78% mAP, surpassing all competing methods by a significant margin. It also achieves the highest mAP under the GG and AG settings (87.62% and 88.62%, respectively), while maintaining highly competitive Rank-1 scores across the remaining protocols. Compared with the previous best AG-ReIDv2 [16], GSAlign improves the average Rank-1 accuracy by +2.7% and the mAP by +6.5%, highlighting its robust cross-domain generalization ability.
Overall, these results consistently verify that GSAlign not only achieves superior retrieval accuracy but also exhibits remarkable robustness across diverse aerial-ground configurations. The consistent gains over both synthetic (CARGO) and real-world (AG-ReID and AG-ReID v2) datasets validate the effectiveness and strong generalization of our geometric and semantic alignment design.
@@
| BoT [41] | 85.40 | 77.03 | 84.65 | 75.90 | 89.77 | 80.48 | 84.65 | 76.90 |
| Explain [6] | 87.70 | 79.00 | 87.35 | 78.24 | 93.67 | 83.14 | 87.73 | 79.08 |
| VDT [8] | 86.46 | 79.13 | 86.14 | 78.12 | 90.00 | 82.21 | 85.26 | 78.52 |
| AG-ReIDv2 [16] | 88.77 | 80.72 | 87.86 | 78.51 | 93.62 | 84.85 | 88.61 | 80.11 |
| SeCap [49] | 88.12 | 80.84 | 88.24 | 79.99 | 91.44 | 84.01 | 87.56 | 80.15 |
| GSAlign | 91.47 | 89.78 | 88.29 | 87.62 | 93.30 | 91.84 | 88.12 | 88.62 |

### v2

@@ Abstract
Aerial-Ground person re-identification (AG-ReID) is an emerging yet challenging task that aims to match pedestrian images captured from drastically different viewpoints, typically from unmanned aerial vehicles (UAVs) and ground-based surveillance cameras. The task poses significant challenges due to extreme viewpoint discrepancies, occlusions, and domain gaps between aerial and ground imagery. While prior works have made progress by learning cross-view representations, they remain limited in handling severe pose variations and spatial misalignment. To address these issues, we propose a Geometric and Semantic Alignment Network (GSAlign) tailored for AG-ReID. GSAlign introduces two key components to jointly tackle geometric distortion and semantic misalignment in aerial-ground matching: a Learnable Thin Plate Spline (LTPS) Module and a Dynamic Alignment Module (DAM). The LTPS module adaptively warps pedestrian features based on a set of learned keypoints, effectively compensating for geometric variations caused by extreme viewpoint changes. In parallel, the DAM estimates visibility-aware representation masks that highlight visible body regions at the semantic level, thereby alleviating the negative impact of occlusions and partial observations in cross-view correspondence. A comprehensive evaluation on CARGO with four matching protocols demonstrates the effectiveness of GSAlign, achieving significant improvements of +18.8% in mAP and +16.8% in Rank-1 accuracy over previous state-of-the-art methods on the aerial-ground setting.
@@ 1 Introduction
Person re-identification (ReID), aiming to address the problem of matching people over a distributed set of nonoverlapping cameras, has attracted intensive attention in the last few years due to its wide applications in surveillance systems. Although traditional person re-identification (ReID) has achieved remarkable progress He et al. (2021); Li et al. (2023b); Tan et al. (2024a); Zhang et al. (2023); Li et al. (2023a), there is growing interest in aerial-ground ReID (AG-ReID) settings, driven by the increasing deployment of unmanned aerial vehicles (UAVs) and low-altitude platforms in surveillance applications Nguyen et al. (2023a); Zhang et al. (2020, 2024b); Chen et al. (2023); Cao et al. (2024). This, in practice, puts the Re-ID problem in an aerial-ground setting and requires the approaches to properly handle both the significant geometric view-variation and semantic misalignment.
Despite its potential, AG-ReID remains a highly underexplored and technically challenging problem. Unlike conventional ReID tasks Tan et al. (2024c); Li et al. (2021b); Wang et al. (2025a); Feng et al. ; Jia et al. (2025), AG-ReID suffers from extreme viewpoint disparities, where aerial images exhibit severe top-down perspectives while ground views contain mostly frontal or profile views. These differences cause significant geometric distortions and drastic appearance changes, which conventional ReID models that are primarily trained on ground-view data are ill-equipped to handle. In addition, AG-ReID often involves frequent occlusions and unbalanced visibility across views, further complicating cross-view matching.
To mitigate these challenges, several recent studies have attempted to adapt existing ReID techniques to the aerial-ground scenario Nguyen et al. (2024); Zhang et al. (2024b). These methods seek to learn modality-specific representations or map features into a shared embedding space using metric learning or adversarial objectives. Others employ cross-view decomposition strategies or design view-invariant constraints to bridge the modality gap. However, these methods tend to focus on global alignment and overlook two critical issues: (1) the severe geometric distortion induced by cross-view perspectives, and (2) semantic misalignment caused by partial occlusions and varying visible body regions. These limitations lead to suboptimal performance, especially under large viewpoint gaps.
@@ 2 Related Work
Building upon well-established ground surveillance infrastructure, advanced person re-identification algorithms have made significant progress across a range of scenarios, including general scenarios Sun et al. (2019); Chen et al. (2019); He et al. (2021); Zhang et al. (2023), occluded scenarios Wang et al. (2020); Li et al. (2021b); Tan et al. (2024b); Ye et al. (2024), cross-modal scenarios Shao et al. (2023); Qin et al. (2024); Zhao et al. (2024); Luo et al. (2025), multi-spectral scenarios Wang et al. (2022); Zhang et al. (2024a); Wang et al. (2025d), and cross-spectral scenarios Kim et al. (2023); Tan et al. (2023); Zhang et al. (2025b); Zhang and Wang (2023); Ren and Zhang (2024).
However, with the proliferation of unmanned aerial vehicles (UAVs) and low-altitude platforms augmenting traditional surveillance systems, person ReID research is increasingly extending to aerial-view scenarios Nguyen et al. (2023b). Initial efforts in this direction focused on aerial-aerial retrieval tasks using datasets such as UAV-Human Li et al. (2021a) and PRAI-1581 Zhang et al. (2020). More recently, attention has shifted towards connecting the aerial views with traditional ground systems. Aerial-Ground person ReID (AG-ReID), where the query and gallery images originate from fundamentally different viewpoints, typically aerial and ground-level perspectives. This setup introduces severe spatial and semantic discrepancies, often leading to substantial degradation in identity-preserving cues such as pose, silhouette, and clothing texture Schumann and Metzler (2017). Several approaches have been proposed to mitigate these challenges. Nguyen et al. Nguyen et al. (2023a) proposed a dual-stream structure guided by attributes to enhance semantic disentanglement, later extended to a three-stream model with modality-aware supervision Nguyen et al. (2024). Zhang et al. Zhang et al. (2024b) developed the View-Decoupled Transformer (VDT), which explicitly models and separates viewpoint-specific and viewpoint-invariant features. SD-ReID Wang et al. (2025c) proposes a generative framework that leverages diffusion models to synthesize view-specific features and enhance view-invariant person representations. In addition, LATex Zhang et al. (2025a) incorporates attribute-based text knowledge via prompt-tuning strategies on vision-language models, explicitly exploiting the viewpoint-invariant nature of person attributes to improve cross-view retrieval. Despite these advances, existing AG-ReID methods still face difficulties in handling extreme spatial distortions and in leveraging visible semantic regions effectively. To overcome these limitations, we propose GSAlign, a novel framework that performs instance-adaptive geometric transformation and visibility-aware semantic alignment, promoting robust and spatially consistent feature learning across aerial and ground domains.
@@ 3.1 Overview
The overall framework of the Geometric and Semantic Alignment Network (GSAlign), illustrated in Fig. 2, is built upon a ViT-Base backbone and designed to address feature misalignment under complex multi-view conditions. Inspired by the success of the View-Decoupled Transformer (VDT) Dosovitskiy et al. (2020), we adopt the VDT architecture as the foundation of GSAlign, which effectively captures both global semantics and fine-grained structural cues. Following its design, we introduce an additional view token to model viewpoint-specific information. The token sequence during training is defined as:
@@ 3.3 Dynamic Alignment Module
In person re-identification, training images often contain occlusions, background noise, or identity-irrelevant regions. Directly encoding features from the entire image may lead the model to learn redundant or non-discriminative representations. To address this issue, we explore whether the visibility patterns of a given input image can be used to dynamically modulate the feature representations of other samples. Inspired by DPM Tan et al. (2022), we propose the Dynamic Alignment Module (DAM), which leverages input-guided masks to improve semantic alignment. Specifically, DAM treats the mean feature representation of images with the same identity as its ideal prototype, assuming that noise or occlusion in different samples manifests as partial loss of this prototype’s semantic structure. Based on this formulation, DAM introduces a dynamic channel-wise mask generator based on the input and modulates prototypes by referencing the input-aware mask. This allows the model to suppress irrelevant or noisy dimensions and emphasize input-relevant subspaces in a content-aware manner.
@@ 4.1 Implementation
Datasets and Evaluation Protocols. We conduct experiments on the CARGO dataset Zhang et al. (2024b), a large-scale benchmark specifically designed for aerial-ground person re-identification (AG-ReID). CARGO contains 108,563 images of 5,000 synthetic identities captured by 13 cameras (5 aerial and 8 ground), with diverse conditions including extreme viewpoint changes, resolution variations, illumination shifts, and occlusions. The dataset is constructed in a simulated urban environment using MakeHuman for identity modeling and Unity3D for camera deployment.
Following the standard protocol, we use 51,451 images with 2500 identities for training and 51,024 images with the remaining 2,500 identities for testing. The testing phase contains four evaluation protocols: ALL, ground to ground( GG), aerial to aerial ( AA), and aerial to ground (AG), where each protocol focuses on a specific matching scenario. To further verify the generalization ability of GSAlign in real-world scenarios, we also conduct comparative experiments on the AG-ReID Nguyen et al. (2023a) and AG-ReID v2 Nguyen et al. (2024) datasets, which contain real aerial and ground imagery captured in outdoor environments. We use Cumulative Matching Characteristics (CMC) at Rank-1 accuracy, mean Average Precision (mAP), and mean Inverse Negative Penalty (mINP) Ye et al. (2021) as evaluation metrics.
Implementation Details. All implementation settings strictly follow the protocol of VDT Zhang et al. (2024b) for fair comparison. We adopt ViT-Base as the backbone, initialized with ImageNet-21K pretraining. Input images are resized to , and horizontal flipping is applied during training for data augmentation. The model is optimized using AdamW with weight decay of 0.05 and a cosine learning rate schedule. The initial learning rate is set to with linear warm-up for the first 20 epochs. is set to 0.1 during the training. We train for 120 epochs in total, using a batch size of 64.
@@ 4.2 Comparison with State-of-the-Arts
| SBS He et al. (2023) | 50.32 | 43.09 | 29.76 | 72.31 | 62.99 | 48.24 | 67.50 | 49.73 | 29.32 | 31.25 | 29.00 | 18.71 |
| PCB Sun et al. (2019) | 51.00 | 44.50 | 32.20 | 74.10 | 67.60 | 55.10 | 55.00 | 44.60 | 27.00 | 34.40 | 30.40 | 20.10 |
| BoT Luo et al. (2019) | 54.81 | 46.49 | 32.40 | 77.68 | 66.47 | 51.34 | 65.00 | 49.79 | 29.82 | 36.25 | 32.56 | 21.46 |
| MGN Wang et al. (2018) | 54.81 | 49.08 | 36.52 | 83.93 | 71.05 | 55.20 | 65.00 | 52.96 | 36.78 | 31.87 | 33.47 | 24.64 |
| VV Kuma et al. (2019); Kumar et al. (2020) | 45.83 | 38.84 | 39.57 | 72.31 | 62.99 | 48.24 | 67.50 | 49.73 | 29.32 | 31.25 | 29.00 | 18.71 |
| AGW Ye et al. (2021) | 60.26 | 53.44 | 40.22 | 81.25 | 71.66 | 58.09 | 67.50 | 56.48 | 40.40 | 43.57 | 40.90 | 29.39 |
| BAU Cho et al. (2024) | 45.20 | 38.40 | - | 61.60 | 51.20 | - | 50.00 | 42.60 | - | 40.40 | 36.70 | - |
| PAT Ni et al. (2023) | 37.90 | 15.30 | - | 52.70 | 24.20 | - | 50.00 | 23.10 | - | 35.10 | 15.50 | - |
| DTST Wang and Pishgar (2024) | 64.42 | 55.73 | 41.92 | 78.57 | 72.40 | 62.10 | 80.00 | 63.31 | 44.67 | 50.53 | 43.49 | 29.46 |
| ViT Dosovitskiy et al. (2020) | 61.54 | 53.54 | 39.62 | 82.14 | 71.34 | 57.55 | 80.00 | 64.47 | 47.07 | 43.13 | 40.11 | 28.20 |
| VDT Zhang et al. (2024b) | 64.10 | 55.20 | 41.13 | 82.14 | 71.59 | 58.39 | 82.50 | 66.83 | 50.22 | 48.12 | 42.76 | 29.95 |
@@
In this section, we compare our proposed GSAlign with a range of state-of-the-art person ReID methods, including both conventional CNN-based models (e.g., PCB Sun et al. (2019), BoT Luo et al. (2019), MGN Wang et al. (2018)) and Transformer-based frameworks (e.g., ViT Dosovitskiy et al. (2020), VDT Zhang et al. (2024b)). As shown in Table 1, GSAlign achieves the best overall performance across all four protocols. Specifically, under the most comprehensive ALL protocol, GSAlign achieves 65.06% Rank-1, 57.95% mAP, and 44.97% mINP, outperforming the strong baseline VDT Zhang et al. (2024b) by +0.96%, +2.75%, and +3.84%, respectively. On the most challenging AG cross-view setting, our method surpasses all previous methods by a large margin, reaching 64.89% Rank-1, 61.55% mAP, and 52.81% mINP, significantly ahead of the second-best VDT by +16.77%, +18.79%, and +22.86%, respectively. Although GSAlign is not specifically designed for single-view matching protocols, it does not introduce any negative effects in such settings. On both AA and GG, GSAlign still achieves competitive performance. We attribute this to the fact that, in single-view scenarios, the features are already well aligned due to the limited viewpoint variation, thus requiring minimal additional geometric or semantic correction. In conclusion, these results clearly demonstrate that our proposed GSAlign not only achieves strong global retrieval accuracy but also exhibits exceptional robustness in the presence of extreme viewpoint changes, validating the effectiveness of our geometric and semantic alignment design.
@@
| ViT Dosovitskiy et al. (2020) | 81.47 | 72.61 | - | 82.85 | 73.39 | - |
| VDT Zhang et al. (2024b) | 82.91 | 74.44 | 51.06 | 86.59 | 78.57 | 52.87 |
| GSAlign | 83.75 | 75.01 | 52.11 | 84.10 | 77.73 | 53.63 |
Furthermore, we evaluate the proposed GSAlign on the real-world AG-ReID Nguyen et al. (2023a) and AG-ReID v2 Nguyen et al. (2024) datasets, and compare it with a wide range of state-of-the-art ReID methods, as summarized in Table 2 and Table 3. On the AG-ReID dataset, GSAlign achieves strong and consistent performance under both cross-view protocols. Under the AG setting, GSAlign obtains the best results across all metrics, reaching 83.75% Rank-1, 75.01% mAP, and 52.11% mINP, outperforming the view-aware baseline VDT by +0.84%, +0.57%, and +1.05%, respectively. Under the inverse GA protocol, GSAlign achieves competitive Rank-1 (84.10%) and mAP (77.73%) scores, while yielding the highest mINP of 53.63%, demonstrating its effectiveness in hard-sample and long-tail retrieval under severe viewpoint changes. On the more challenging AG-ReID v2 dataset, which features more realistic and diverse scenes, GSAlign continues to exhibit strong generalization ability. As shown in Table 3, GSAlign achieves the highest mAP across all four protocols, including AG, GA, AW, and WA, with mAP values of 81.38%, 81.05%, 83.98%, and 80.90%, respectively. Although some methods attain slightly higher Rank-1 accuracy under specific settings, GSAlign remains highly competitive while delivering more stable overall retrieval performance.
In summary, the experimental results on both AG-ReID and AG-ReID v2 datasets demonstrate that GSAlign effectively enhances cross-view feature alignment and retrieval robustness. The consistent gains in mAP and mINP across diverse protocols validate the effectiveness and robustness of our geometric and semantic alignment design in real-world aerial–ground ReID scenarios.
@@
| BoT Luo et al. (2019) | 85.40 | 77.03 | 84.65 | 75.90 | 89.77 | 80.48 | 84.65 | 76.90 |
| Explain Nguyen et al. (2023a) | 87.70 | 79.00 | 87.35 | 78.24 | 93.67 | 83.14 | 87.73 | 79.08 |
| AG-ReIDv2 Nguyen et al. (2024) | 88.77 | 80.72 | 87.86 | 78.51 | 93.62 | 84.85 | 88.61 | 80.11 |
| SeCap Wang et al. (2025b) | 88.12 | 80.84 | 88.24 | 79.99 | 91.44 | 84.01 | 87.56 | 80.15 |
| VDT Zhang et al. (2024b) | 86.46 | 79.13 | 86.14 | 78.12 | 90.00 | 82.21 | 85.26 | 78.52 |
| GSAlign | 87.86 | 81.38 | 88.02 | 81.05 | 90.63 | 83.98 | 87.31 | 80.90 |